1Mg= ___________μg [exponent for M ____, exponent for μ ____] 



1
Expert's answer
2021-01-18T01:47:11-0500

1Mg = 10-3g.

1μg = 10-6 g.


So , 1 / 10-3 Mg = 1 / 10-6 μg.

So , 103Mg= 106μg.

So , 1Mg = 103μg.

So , 1 Mg = 103μg. ANSWer .